Let the charge q and -q be placed alternatively along the x-axis at position x=2m,x=3m,x=5m and x=8 m.what is the electric field at x=0 due to these changes?

1
Expert's answer
2022-06-10T18:32:38-0400

"E=kq(\frac 1{x_1^2}-\frac 1{x_2^2}+\frac 1{x_3^2}-\frac 1{x_4^2})."